Find four consecutive odd integers such that the sum of the second and the fourth is 63 more than one-fifth of the third.....that's my question...can you show the solution?????

Answer by Edwin McCravy(20062)


Find four consecutive odd integers such that the sum of the second 
and the fourth is 63 more than one-fifth of the third.....that's my
question...can you show the solution?????

"Four consecutive odd integers" refers to the set {1,3,5,7} or
the set {3,5,7,9} or the set {9,11,13,15} or the set {77,79,81,83} or
any set of four such odd numbers in a row.  

To work such problems you start out by

1. letting x represent the smallest of four consecutive odd integers.

2. Then x+2 represents the second of four consecutive odd integers,

3. Then x+4 represents the third of four consecutive odd integers.

4. Then x+6 represents the fourth of four consecutive odd integers

>>...the sum of the second and the fourth is 63 more than one-fifth 
of the third...<<

Replace the words "the sum of the second and the fourth" by
(x+2) + (x+6)

So we have

>>...(x+2) + (x+6) is 63 more than one-fifth of the third...<<
  
Replace the words "one-fifth of the third" by "(1/5)(x+4)"  

So we now have

>>...(x+2) + (x+6) is 63 more than (1/5)(x+4)...<< 

To get 63 more than (1/5)(x+4) we add 63 to it, so replace
the words "63 more than (1/5)(x+4)" by "(1/5)(x+4) + 63" 

So now we have

>>...(x+2) + (x+6) is (1/5)(x+4) + 63...<< 

Finally replace the word "is" by "=" and we have completely
gotten rid of the English, and all that's left is algebra:

(x+2) + (x+6) = (1/5)(x+4) + 63

To solve that first multiply every term by LCD = 5. That is,
put 5 as a multiplier in front of every term on both sides

5(x+2) + 5(x+6) = 5(1/5)(x+4) + 5(63)

Replace the 5(1/5) on the right by 1, and the 5(63) by 315

5(x+2) + 5(x+6) = 1(x+4) + 315

Now distribute to remove parentheses, combine like signs,
solve, and you get

x = 31

So since the four cnsecutive integers are x, x+2, x+4 and x+6,
are respectively 31, 33, 35 and 37

Edwin